Consider this reaction:

N2(g) + 3H2(g) <==> 2NH3(g)

If the Kc for the reaction is 4.5 x 10^4 at 127 degrees celsius, what is the equilibrium constant, Kp at the same temperature?
Answer #1

Date ....... Na (9) + 3 H2 (g) ke= 4.5*104 → 2NH3 (9) at 127°C osing the relation between kp & ke l.e . kp= kc (R7) oh ke= 4.